OpenStudy (anonymous):

SOLVE _____ 3 _____ [√ 3^3x ] = √ 3^5x * 27^4x-1

OpenStudy (anonymous):

what is that 3 on top? is that an exponent? _____ 3 <--THIS _____ [√ 3^3x ] = √ 3^5x * 27^4x-1

jhonyy9 (jhonyy9):

here is chaos

OpenStudy (anonymous):

This doesn't look so terrible if you just write everything with exponents and simplify. eg the lhs is 3^(9x/2).

OpenStudy (anonymous):

math.yes it is the exponent

OpenStudy (anonymous):

\[\sqrt{3^{3x}}^3 = \sqrt{3^{5x}}∗27^{4x-1}\] This is how I'm reading it